# Learning Mamba as a Continual Learner

### v1

@@ 4.4 Ablation Studies
Hyper-parameter of selectivity regularization loss. We conducted an ablation study to assess the influence of training loss hyper-parameter on our Mamba model’s efficacy. Specifically, this study involved adjusting the $\lambda$ values within our selectivity regularization loss, experimenting with hyper-parameters set at $\{0.1,0.2,0.5,1.0,2.0\}$, as depicted in Fig. 5. The results indicate that these variations have a minor impact on our Mamba model’s performance. Consequently, we selected a $\lambda$ value of $0.5$ for our experiments.
SSM state size. In Fig. 5, we evaluate the impact of varying the SSM state size on the performance of our methods. We conducted experiments on ImageNet-1K and Cifar-100, training MambaCL with state sizes of 16, 32, 64, 128, and 256. The results show consistent performance improvement as state size increases. To balance performance and computational cost, we set the state size as 128.
Figure 4: Ablation of varying $\lambda$ in training loss.
Figure 5: Ablation of varying SSM state size.
Table 6: Different Mamba architectures on 20-tasks 5-shot MCL.
@@
Different architectures. In Table 6, we present an ablation study comparing different Mamba architectures in our MambaCL, including Mamba-1, MambaFormer (Park et al., 2024), and Mamba-2. MambaFormer is a hybrid model that integrates the vanilla attention mechanism of Mamba-1 and replaces the transformer’s positional encoding with a Mamba block. The results in Table 6 demonstrate that MambaFormer also achieved performance comparable to that of the transformer. However, Mamba-2 performed better on Cifar-100 than the other variants.
@@
Mamba+MoE. In Table 6, we present experiments where Mamba was enhanced with Mixture of Experts (MoE), incorporating twelve 2-layer MLP expert networks with a dense-MoE router following each Mamba Block, resulting in improved performance. Additionally, we include performance metrics for vanilla and linear transformers for reference.
Computational cost. In Table 7, we detail various aspects of computational cost using our implementation in PyTorch [27], executed on an NVIDIA 4090 GPU and an INTEL I9-14900k CPU. We specifically report the costs associated with meta-testing at a batch size of 1. Notably, Mamba, characterized by fewer parameters and increased processing speed, achieves performance that either matches or surpasses that of the vanilla transformer.
@@ 5 Conclusion
In this paper, we tried to answer a question – Can attention-free Mamba perform well on MCL? We formulate the SSM and Mamba as a sequence prediction-based continual learner and meta-learn it on CL episodes. A selectivity regularization is introduced for meta-learning the models. Comprehensive experiments show that Mamba performs well across diverse MCL scenarios, significantly outperforming attention-free methods and matching or exceeding Transformers’ performance with fewer parameters and computations. In challenging scenarios with global structures, domain shifts, and long sequences, Mamba demonstrates obvious reliability, generalization, and robustness.
Limitations and future work. This study can be extended to larger-scale datasets and offline CL settings. Beyond the current MCL framework, we aim to explore the online meta-continual learning paradigm to broaden the applicability of our approach to a wider range of scenarios.
@@ A.1 General Image Classification Tasks
Cifar-100 (Krizhevsky & Hinton, 2009) dataset consists of 60,000 images across 100 classes, each with 600 images. We select 60 classes at random for meta-training and use the remaining 40 for meta-testing.
ImageNet-1K (Russakovsky et al., 2015) dataset comprises over one million labeled images distributed across 1,000 categories. We select 600 classes at random for meta-training and use the remaining 400 for meta-testing.
ImageNet-R(endition) (Russakovsky et al., 2015) extends 200 ImageNet classes with a compilation of 30,000 images tailored for robustness research.
Celeb (Guo et al., 2016) is a large-scale facial image collection featuring approximately 10 million images of 100,000 celebrities. We randomly allocated 1,000 classes for meta-testing and assigned the remaining classes to meta-training.
Casia Chinese handwriting (Liu et al., 2011) dataset encompasses a total of 7,356 character classes with 3.9 million images. We randomly selected 1,000 classes for the meta-testing and allocated the remaining classes for meta-training.
Omniglot (Lake et al., 2015) is a collection of 1,632 handwritten characters from 50 different alphabets. The meta-training set comprises 963 classes, while the meta-testing set includes 660 classes, with each class containing 20 images.
@@ A.2 Fine-grained Recognition Tasks
CUB-200-2011 (Wah et al., 2011) is a widely used fine-grained visual categorization dataset, comprising 11,788 images across 200 bird subcategories. We randomly selected 80 classes for the meta-testing and allocated the remaining classes for meta-training.
Stanford Dogs (Khosla et al., 2011) dataset comprises 20,580 images spanning 120 global dog breeds, divided into 12,000 training images and 8,580 testing images. We select 48 classes at random for meta-testing and use the remaining 72 for meta-training.
Stanford Cars (Krause et al., 2013) comprises 16,185 images across 196 car classes, primarily captured from the rear perspective. We select 80 classes at random for meta-testing and use the remaining 40 for meta-training.
FGVC-Aircraft (Maji et al., 2013) dataset comprises 10,200 images across 102 aircraft model variants, each represented by 100 images, primarily consisting of airplanes. We randomly selected 40 classes for the meta-testing and allocated the remaining classes for meta-training.
@@ A.3 Large Domain Shift Tasks
DomainNet (Peng et al., 2019) dataset is a benchmark for domain adaptation, encompassing common objects organized into 345 classes across six domains: clipart, real, sketch, infograph, painting, and quickdraw. We evaluate model adaptability to out-of-domain data by using one domain for meta-testing and the remaining domains for meta-training.
@@ A.3.1 Regression Tasks
Appendix B Additional experimental Details
Table 8 presents the configurations of the models employed in our experiments.
@@ C.1 Effects of Selectivity Regularization and Meta-Training Loss Curves
C.2 More Ablation Studies on Regularization Strength
In Fig. 5, we conducted ablation study to assess the influence of regularization strengths on our Mamba’s efficacy. Fig. 7 illustrates more ablation studies assessing the impact of regularization strengths $\lambda$ by setting is as $\{0.1,0.2,0.5,1.0,2.0\}$, across multiple models on both ImageNet-1K and Cifar-100 datasets. The results demonstrate that all models exhibit stability within a wide and appropriate range of $\lambda$, providing evidence of consistent patterns. In our experiments, without losing generality, all models employed a regularization strength of 0.5 by default.
C.3 More Ablation Studies on Learning Rates
@@ C.4 Additional Generalization Analyses
Without the regularization, models struggle to converge and exhibit significant oscillations during training, as shown in Fig. 6. In Sec. 4.2 and Fig. 3, we conducted generalization analyses of various models by conducting meta-testing on the episodes different from the meta-training settings. Specifically, we apply the models meta-trained with 20-task-5-shot episodes on the meta-testing episodes with varying numbers of tasks or shots or the episodes contaminated by noise. The results show that Mamba shows better generalization ability to unseen scenarios and Transformer shows more meta-overfitting issues. To validate that the results are not relevant to the regularization, we evaluated various models with a small regularization strength ($\lambda=0.1$) to assess the impact of regularization on this generalization experiment and the meta-overfitting issue. The results indicate that regularization strengths of 0.1 (Fig. 9) and 0.5 (Fig. 3) lead to similar phenomena across different models.
@@ Appendix D Visualization of Attention and Selectivity Pattern
Given meta-learned sequence models as the continual learner, the models process the samples in sequence in the meta-test CL process. To analyze the behaviors of these models, we visualize the attention weights of Transformers and the associative weights of Mamba (as discussed in Sec. 3.2.2) to demonstrate their attention and selectivity patterns, respectively. In a meta-testing episode, given a trained model and a sequence of samples, the prediction for a given ${\mathbf{x}}^{\text{test}}$ is produced based on the attention or implicit association of seen samples in the sequence. Visualizing the attention and selectivity patterns can empirically show how the models make predictions. For the standard benchmarking case, Fig. 10 shows that both Transformer and Mamba can effectively associate seen samples with query inputs, leading to the results as shown in Table 2.
Specifically, we use this visualization to analyze how different models perform in the generalization studies (discussed in Sec. 4.2), i.e., generalizing to meta-testing cases that are different from meta-training cases.
@@ D.1 Visualization Analyses for Generalization to Different Stream Length
The experiments shown in Fig. 3 and Fig. 3 validate the generalization ability of models by meta-testing on CL episodes/sequences that differ from those seen during meta-training. Specifically, the models are meta-trained on 20-task, 5-shot MCL episodes and meta-tested on episodes with task and shot numbers exceeding those in meta-training. Transformers generally converge more easily during meta-training compared to Mamba, due to their strong fitting ability. However, this advantage may also lead to meta-overfitting.
To analyze how different models perform on these sequences, we visualize the final layer attention weights of Transformer and the corresponding selective scores (associative indicators) of Mamba, between various test shots (queries) and a single MCL train episode (prompt) of both Mamba and Transformer. Note that Mamba does not have explicit attention weights, we compute the scores relying on the connection between Mamba and Transformers described in Sec. 3.2.2. Specifically, we computed the parameters ${\mathbf{C}}_{\text{test}}$ and ${\mathbf{B}}$ (${\mathbf{C}}_{\text{test}}{\mathbf{B}}^{\top}$) within its SSMs to compare its behavior with the attention matrix (${\mathbf{Q}}_{\text{test}}{\mathbf{K}}^{\top}$) of Transformers, where ${\mathbf{C}}_{\text{test}}\in{\mathbb{R}}^{1\times C}$ and ${\mathbf{Q}}_{\text{test}}\in{\mathbb{R}}^{1\times C}$ correspond to the row of the test shot. Both models are meta-trained on a 20-task, 5-shot setting using the ImageNet-1K dataset.
@@ D.2 Visualization Analysis of Generalization to Noise-Contaminated Episodes
In the experiments, the modes are meta-trained on noise-free episodes. And the noise is added on randomly selected samples/shots in the meta-testing episodes. The task can also be seen as validating the ability of ignoring the irrelevant samples or contaminated outlier samples in the sequences. To directly show how the models work in this scenarios, we visualized the final layer attention weights for test shots compared to training shots for both Mamba and Transformer, each meta-trained in a 20-task, 5-shot setting. During meta-testing, these models processed a 20-task, 5-shot episode with five noisy input shots (shot index: 8, 18, 39, 61, 75) at noise strengths of 1 (Fig. 13), 2 (Fig. 14), and 6 (Fig. 15). The results indicate that the Transformer meta-trained on clean episodes tend to produce extreme attention weights (either very high or very low) on noisy or outlier shots, whereas Mamba is less affected. This observation suggests that Transformers’ learned attention mechanisms tend to associate samples based on local and independent representations. In contrast, Mamba performs more effectively by selectively associating relevant information and leveraging its recurrently updated latent state, which accumulates global sequence information.

### v4

@@ 4.3 Ablation Studies
Table 6: Different Mamba architectures on 20-task 5-shot MCL.
@@
Hyper-parameter of selectivity regularization loss. We conducted ablation studies to assess the impact of the selectivity regularization weight $\lambda$ on our Mamba model’s performance, varying it across ${0.1,0.2,0.5,1.0,2.0}$ (Fig. 5). The results indicate that performance remains largely stable across different $\lambda$ values, suggesting Mamba’s robustness to this hyper-parameter. Based on these findings, we set $\lambda=0.5$ for all reported experiments.
SSM state size. In Fig. 5, we assess the impact of varying SSM state size on the performance of our methods. We conducted experiments on ImageNet-1K and Cifar-100, training the MambaCL with state sizes of 16, 32, 64, 128, and 256. The results demonstrate a consistent improvement in performance as the state size increases. To optimally balance performance with computational efficiency, we selected a state size of 128 for our experiments.
Different architectures. Table 6 presents an ablation study comparing Mamba-1, MambaFormer [45], and Mamba-2 within our MambaCL framework. MambaFormer is a hybrid model that integrates vanilla attention into Mamba and replaces Transformer positional encoding with a Mamba block. The results show MambaFormer performs comparably to Transformers, while Mamba-2 achieves the best performance on the Cifar-100 dataset.
Mamba+MoE. In Table 6, we present experiments in which Mamba is augmented with mixture-of-experts (MoE), incorporating twelve 2-layer MLP expert networks with a dense-MoE router following each Mamba Block, resulting in improved performance. Additionally, we include performance metrics for vanilla and linear transformers as references.
@@
Computational cost. In Table 7, we detail various aspects of computational cost using our implementation in PyTorch, executed on an NVIDIA 4090 and an INTEL I9-14900K. We specifically report the costs associated with meta-testing at a batch size of 1. Mamba achieves comparable or superior performance to the vanilla Transformer, while being more efficient in parameters and speed.
@@ 5 Conclusion
In this paper, we try to answer the question – Can attention-free Mamba perform well on MCL? We formulate SSMs and Mamba as a sequence prediction-based continual learner and meta-learn it on CL episodes. A selectivity regularization is introduced for meta-continual learning the models. Comprehensive experiments show that Mamba performs well across diverse MCL scenarios, significantly outperforming other attention-free methods and matching or exceeding Transformers’ performance with fewer parameters and computations. In challenging scenarios with global structures, domain shifts, and long sequences, Mamba demonstrates obvious reliability, generalization, and robustness.
Limitations and future work. This study can be extended to larger-scale datasets and offline CL settings. Future work also includes exploring other types efficient models, such as efficient Transformers with compressed key-value caches and efficient CL for large-scale foundation models. Beyond the current MCL framework, we aim to expand its applicability to a broader range of scenarios and models.
@@ A.1 General Image Classification Tasks
Cifar-100 [30] dataset consists of 60,000 images across 100 classes, each with 600 images. We select 60 classes at random for meta-training and use the remaining 40 for meta-testing.
ImageNet-1K [53] dataset comprises over one million labeled images distributed across 1,000 categories. We select 600 classes at random for meta-training and use the remaining 400 for meta-testing.
ImageNet-R [53] extends 200 ImageNet classes with a compilation of 30,000 images tailored for robustness research.
Celeb [19] is a large-scale facial image collection featuring approximately 10 million images of 100,000 celebrities. We randomly allocated 1,000 classes for meta-testing and assigned the remaining classes to meta-training.
Casia Chinese handwriting [36] dataset encompasses a total of 7,356 character classes with 3.9 million images. We randomly selected 1,000 classes for the meta-testing and allocated the remaining classes for meta-training.
Omniglot [31] is a collection of 1,632 handwritten characters from 50 different alphabets. The meta-training set comprises 963 classes, while the meta-testing set includes 660 classes, with each class containing 20 images.
@@ A.2 Fine-grained Recognition Tasks
CUB-200-2011 [62] is a widely used fine-grained visual categorization dataset, comprising 11,788 images across 200 bird subcategories. We randomly selected 80 classes for the meta-testing and allocated the remaining classes for meta-training.
Stanford Dogs [10] dataset comprises 20,580 images spanning 120 global dog breeds, divided into 12,000 training images and 8,580 testing images. We select 48 classes at random for meta-testing and use the remaining 72 for meta-training.
Stanford Cars [29] comprises 16,185 images across 196 car classes, primarily captured from the rear perspective. We select 80 classes at random for meta-testing and use the remaining 40 for meta-training.
FGVC-Aircraft [40] dataset comprises 10,200 images across 102 aircraft model variants, each represented by 100 images, primarily consisting of airplanes. We randomly selected 40 classes for the meta-testing and allocated the remaining classes for meta-training.
@@ A.3 Large Domain Shift Tasks
DomainNet [46] dataset is a benchmark for domain adaptation, encompassing common objects organized into 345 classes across six domains: clipart, real, sketch, infograph, painting, and quickdraw. We evaluate model adaptability to out-of-domain data by using one domain for meta-testing and the remaining domains for meta-training.
@@ A.3.1 Regression Tasks
Table 8: Model Configurations
@@
Appendix B Implementation Details
We conduct our main experiments on a single NVIDIA A100 GPU, repeating each experiment five times and reporting the mean and standard deviation. Results are recorded upon convergence on the meta-training set. We use a batch size of 16 and the Adam optimizer, with an initial learning rate of $1\times 10^{-4}$ that decays by $0.5$ every $10,000$ steps. All models follow a consistent setup to ensure fair comparisons, with details provided in the Appendix. For experiments involving training from scratch, we adopt settings from [32] for fairness. For experiments with pre-trained networks, we employ OpenAI/CLIP-ViT-B16 [48, 24] as the image encoder, keeping its parameters frozen and adding a trainable linear projector. Table 8 presents the configurations of the models employed in our experiments.
@@ C.1 Effects of Selectivity Regularization and Meta-Training Loss Curves
C.2 Ablation Studies on Regularization Strength
Fig. 5 in the main paper, we conducted an ablation study to assess the influence of regularization strengths on our Mamba’s efficacy. Fig. 7 illustrates more ablation studies assessing the impact of regularization strengths $\lambda$ by setting is as $\{0.1,0.2,0.5,1.0,2.0\}$, across multiple models on both ImageNet-1K and Cifar-100 datasets. The results demonstrate that all models exhibit stability within a wide and appropriate range of $\lambda$, providing evidence of consistent patterns. In our experiments, without losing generality, all models employed a regularization strength of 0.5 by default.
C.3 Ablation Studies on Learning Rates
@@ C.4 Additional Generalization Analyses
Without the regularization, models struggle to converge and exhibit significant oscillations during training, as shown in Fig. 6. In Sec. 4.2 and Fig. 5 of the main paper, we conducted generalization analyses of various models by conducting meta-testing on the episodes different from the meta-training settings. Specifically, we apply the models meta-trained with 20-task-5-shot episodes on the meta-testing episodes with varying numbers of tasks or shots or the episodes contaminated by noise. The results show that Mamba shows better generalization ability to unseen scenarios and Transformer shows more meta-overfitting issues. To validate that the results are not relevant to the regularization, we evaluated various models with a small regularization strength ($\lambda=0.1$) to assess the impact of regularization on this generalization experiment and the meta-overfitting issue. The results indicate that regularization strengths of 0.1 (9) and 0.5 (5 in the main paper) lead to similar phenomena across different models.
@@ Appendix D Visualization of Attention and Selectivity Pattern
Given meta-learned sequence models as the continual learner, the models process the samples in sequence in the meta-test CL process. To analyze the behaviors of these models, we visualize the attention weights of Transformers and the associative weights of Mamba (as discussed in 3.2.2 in the main paper) to demonstrate their attention and selectivity patterns, respectively. In a meta-testing episode, given a trained model and a sequence of samples, the prediction for a given ${\mathbf{x}}^{\text{test}}$ is produced based on the attention or implicit association of seen samples in the sequence. Visualizing the attention and selectivity patterns can empirically show how the models make predictions. For the standard benchmarking case, Fig. 10 shows that both Transformer and Mamba can effectively associate seen samples with query inputs, leading to the results as shown in Table 2 in the main paper.
Specifically, we use this visualization to analyze how different models perform in the generalization studies (discussed in 4.2 of the main paper), i.e., generalizing to meta-testing cases that are different from meta-training cases.
@@ D.1 Visualization Analyses for Generalization to Different Stream Length
The experiments shown in 5 and 5 in the main paper validate the generalization ability of models by meta-testing on CL episodes/sequences that differ from those seen during meta-training. Specifically, the models are meta-trained on 20-task, 5-shot MCL episodes and meta-tested on episodes with task and shot numbers exceeding those in meta-training. Transformers generally converge more easily during meta-training compared to Mamba, due to their strong fitting ability. However, this advantage may also lead to meta-overfitting.
To analyze how different models perform on these sequences, we visualize the final layer attention weights of Transformer and the corresponding selective scores (associative indicators) of Mamba, between various test shots (queries) and a single MCL train episode (prompt) of both Mamba and Transformer. Note that Mamba does not have explicit attention weights, we compute the scores relying on the connection between Mamba and Transformers described in 3.2.2 in the main paper. Specifically, we computed the parameters ${\mathbf{C}}_{\text{test}}$ and ${\mathbf{B}}$ (${\mathbf{C}}_{\text{test}}{\mathbf{B}}^{\top}$) within its SSMs to compare its behavior with the attention matrix (${\mathbf{Q}}_{\text{test}}{\mathbf{K}}^{\top}$) of Transformers, where ${\mathbf{C}}_{\text{test}}\in{\mathbb{R}}^{1\times C}$ and ${\mathbf{Q}}_{\text{test}}\in{\mathbb{R}}^{1\times C}$ correspond to the row of the test shot. Both models are meta-trained on a 20-task, 5-shot setting using the ImageNet-1K dataset.
@@ D.2 Visualization Analysis of Generalization to Noise-Contaminated Episodes
In the experiments, the modes are meta-trained on noise-free episodes. The noise is added to randomly selected samples/shots in the meta-testing episodes. The task can also be seen as validating the ability to ignore the irrelevant samples or contaminated outlier samples in the sequences. To directly show how the models work in this scenario, we visualized the final layer attention weights for test shots compared to training shots for both Mamba and Transformer, each meta-trained in a 20-task, 5-shot setting. During meta-testing, these models processed a 20-task, 5-shot episode with five noisy input shots (shot index: 8, 18, 39, 61, 75) at noise strengths of 1 (Fig. 13), 2 (Fig. 14), and 6 (Fig. 15). The results indicate that the Transformer meta-trained on clean episodes tends to produce extreme attention weights (either very high or very low) on noisy or outlier shots, whereas Mamba is less affected. This observation suggests that Transformers’ learned attention mechanisms tend to associate samples based on local and independent representations. In contrast, Mamba performs more effectively by selectively associating relevant information and leveraging its recurrently updated latent state, which accumulates global sequence information.
Appendix E Broader impacts
This work presents both promising and cautionary societal implications. On the positive side, MambaCL enables efficient continual learning with low memory and computation costs, making it well-suited for deployment in resource-constrained or privacy-sensitive environments such as mobile devices, robotics, and personalized healthcare systems. Its robustness to domain shifts and noisy inputs further supports real-world reliability. However, our model has not been trained or evaluated on larger-scale datasets or in real-world industry scenarios, so its behavior and potential impacts in such settings remain uncertain. Care should be taken when applying it to more complex or practical environments.